# All-in-one: Understanding and Generation in Multimodal Reasoning with the MAIA Benchmark

### v2

@@ Abstract
We introduce MAIA (Multimodal AI Assessment), a native-Italian benchmark designed for fine-grained investigation of the reasoning abilities of visual language models on videos. MAIA differs from other available video benchmarks for its design, its reasoning categories, the metric it uses, and the language and culture of the videos. MAIA evaluates Vision Language Models (VLMs) on two aligned tasks: a visual statement verification task, and an open-ended visual question-answering task, both on the same set of video-related questions. It considers twelve reasoning categories that aim to disentangle language and vision relations by highlighting the role of the visual input. Thanks to its carefully taught design, it evaluates VLMs’ consistency and visually grounded natural language comprehension and generation simultaneously through an aggregated metric revealing low results that highlight models’ fragility. Last but not least, the video collection has been carefully selected to reflect the Italian culture, and the language data are produced by native-speakers. [^1] Data and code will be available upon acceptance.
@@
Davide Testa1,2, Giovanni Bonetta2, Raffaella Bernardi3, Alessandro Bondielli4,5, Alessandro Lenci5, Alessio Miaschi6, Lucia Passaro4, Bernardo Magnini2 1Università di Roma La Sapienza, 2Fondazione Bruno Kessler (FBK), 3CIMeC, DISI, University of Trento 4Dept. of Computer Science, University of Pisa, 5Dept. of Philology, Literature and Linguistics, University of Pisa 6Istituto di Linguistica Computazionale "A. Zampolli" (CNR-ILC), ItaliaNLP Lab, Pisa davide.testa@uniroma1.it, {dtesta, gbonetta, magnini}@fbk.eu, raffaella.bernardi@unitn.it {name.surname}@unipi.it, alessio.miaschi@ilc.cnr.it
@@ 1 Introduction
Figure 1: Structure of the MAIA benchmark. For each video there are two questions related to 12 reasoning categories. For each question there is a pool of 8 answers, which are associated to a true-false statement pair. The example reported in the Figure is related to Out-of-scope reasoning category.
@@
To address these limitations, we present MAIA: a competence-oriented benchmark consisting of two paired tasks – multiple-choice Visual Statement Verification (VSV) and Open-Ended Visual Question Answering (OEVQA) – containing aligned datapoints. For instance, in the example in Figure 1, for the video showing a pizza cooking in a wood-fired oven, the VSV contains the Truth and False statements (TS and TF) A: There is no cake in the video and B: There is a cake in the video, and the OEVQA contains the (aligned) question What is the cake made of?. Such interleaved data let us evaluate models’ robustness: we evaluate a model positively, only if it performs correctly both on the visual statement verification task (NLU) – it selects There is no cake in the video – and on the Open-ended VQA (NLG) by generating something like There is a pizza, not a cake; while we evaluate it negatively, if it performs correctly only on one of the two tasks. Moreover, the VSV is organized in pools containing 8 TS-FS pairs that differ only on the surface, letting us evaluate the model’s consistency. We categorize questions based on the reasoning they elicit; for instance, the question in Figure 1 is "out-of-scope". MAIA spans twelve reasoning categories, helping highlight the role of language and visual modalities across them. Finally, it implements an all-in-one evaluation philosophy that lets us evaluate both models’ robustness and consistency through its Aggregate Metric. Last but not least, MAIA is based on native Italian videos. To the best of our knowledge, this is the first benchmark for the Italian language on videos.
@@ 2 Related Work
Various types of benchmarks have been proposed since the raise of VLMs. From the single task-oriented benchmarks (e.g. Antol et al. (2015); Das et al. (2017)), attention has now moved to task collections Xu et al. (2024); Lee et al. (2024b) in which models show impressive performance. As in the early phase Johnson et al. (2017); Shekhar et al. (2017); Suhr et al. (2017), such success is mitigated by the use of diagnostic benchmarks, such as Parcalabescu et al. (2022); Thrush et al. (2022); Chen et al. (2023); Bugliarello et al. (2023); Bianchi et al. (2024) and carefully curated benchmarks such as Xiao et al. (2024); Tong et al. (2024). The third type of benchmarks available focus on the VLMs competence in a holistic fashion, evaluating advanced perception and reasoning with domain-specific knowledge Yue et al. (2024b). A similar picture emerges for video-based VLMs. Here as well, early surveys call for careful evaluation (e.g. Zhong et al. (2022)), task-oriented benchmarks show impressive performance Grunde-McLaughlin et al. (2021); Yu et al. (2023), while fine-grained ones pinpoint important weaknesses Kesen et al. (2023), and competence-based analysis highlight there is significant room for improvement in multimodal video understanding (Patraucean et al., 2023). Finally, both Tong et al. (2024) for images and Kesen et al. (2023) for videos manage to highlight VLMs shortcomings by imposing a more stringent task-accuracy metric that account for model consistency across very similar data or correlated competencies. Thanks to the richness of MAIA data collection, we adopt such severe, and hence robust, evaluation code and propose a novel aggregate metric. Building on prior work, MAIA targets a low-resource language and the underexplored video domain. While some benchmarks Das et al. (2024); Zhang et al. (2023) include limited Italian multiple-choice tasks, none focus on high-level reasoning in Italian video contexts or analyze distributional biases in VLMs.
@@
Widely used benchmarks, such as AGQA Grunde-McLaughlin et al. (2021) and MVBench Li et al. (2024c) focus on explicit visual elements (e.g., entity, action, and the spatio-temporal reasoning involving them), instead MAIA’s categories focus on the interplay between language and vision, especially when this relation is implicit or must be inferred, a dimension largely neglected in prior Video QA benchmarks. Yue et al. (2024a) includes multiple-choice and open-ended data points from entirely independent data sets with different origins and content, and reports the average performance across distinct tasks. Instead, MAIA’s NLU and NLG data points are aligned, a unique feature of MAIA framework, as such it introduce an Aggregate Accuracy metric specifically designed to ensure that the model’s performance is evaluated consistently across multiple-choice and open-ended questions derived from the same underlying data. There are few other video-text benchmarks including both these formats (e.g., Zhou et al. (2025); Peng et al. (2024)), however, to the best of our knowledge, none of them attempt to define a unifying metric, as we do in MAIA.
@@ 3 The MAIA Benchmark
Figure 2: Overview of MAIA reasoning categories (we illustrate 9/12 categories, see the Appendix for the examples about the missing ones). For each of the 100 videos, it contains 2 questions for each of the 12 categories; for each question, it has 8 answers, and each of these answers has a corresponding True and False statement pair.
@@ 3.1 Dataset
We outline the steps for collecting, generating, and validating the MAIA dataset. Validation steps consists in a qualitative analysis and revision of the data, when necessary. [^2] Examples are in English for readability. Appendix A provides more details, Italian examples, and translations.
@@
Questions and Answers Collection. The annotation process was carried out in two phases. In the first phase (question creation), $12$ qualified annotators wrote $2$ open-ended questions [^3] Yes/No and audio-based questions were prohibited. per video for each category, ensuring diversity in entities and events. [^4] Annotators were paid €100 for their work. A manual review verified adherence to guidelines and semantic categories. In the second phase (answer collection), we used Prolific [^5] https://www.prolific.com/ to solve the task, targeting Italian-native participants with specific cultural criteria (Appendix A.3). Each annotator answered 12 out of 24 questions per video [^6] Annotators were paid £7 per hour., focusing on detailed, visually grounded responses. Each question was answered by eight annotators, resulting in $19,200$ responses. Two semi-automatic validation checks were applied to the output: ($1$) semantic consistency with the corresponding question, and ($2$) contradiction tests within each answer pool (Appendix A.3) [^7] We found that 90.25% of the 8-answer pools exhibit full agreement, as they do not contain any contradictions. The remaining 9.75% was manually reviewed by an additional annotator to resolve inconsistencies. This shows agreement among annotators, that cannot be measured through standard annotator agreement metrics due to the open-ended nature of the task.. Through a post-processing of the lexicon, we made sure that the final 8-answer pools are lexical diverse. [^8] The lexical overlap within the 8-answer pool is $21.90$%, similarly to the overlap between pairs randomly extracted from each pool ($22.21$%).
@@
Statement Collection. As shown in Figure 2, TSs are descriptive declarative sentences that accurately align with the visual content of videos. TSs describe videos from different semantic perspectives, according to MAIA semantic categories. TSs were generated using GPT-4o (prompts are reported in Figure 12 of the Appendix): for each question, it is given the 8 human generated answers and it is prompted to produce 8 TSs by combining the content of the question with the one of the corresponding answer. Again, post-processing techniques ensured high lexical variability within each pool reducing lexical overlap. FSs are incorrect descriptions automatically generated by prompting GPT-4o and created by modifying elements related to a reasoning category while maintaining the sentence structure (Appendix A.7). FSs were validated through two semi-automatic checks: a check to ensure minimal modification while maintaining incorrectness for the semantic category, and a test to confirm that FSs contradict their corresponding TSs (Appendix A.8). This process produced $19,200$ high-quality FSs aligned with their corresponding TSs.
@@ 3.2 Reasoning Categories
We report the reasoning categories in MAIA. Figure 2 provides examples of aligned question, answer, TS, and FS. More details in Appendix A.2.
@@ 4.1 Task1: Visual Statement Verification
VSV is a multiple-choice task where a model is presented with a true-false statement pair related to a MAIA question (see section 3.1) for a given video, and has to select the true statement. The two statements are randomly assigned to two labels, A and B, and the model is asked to generate only the label. We chose the prompt through an extensive evaluation of $32$ variants ($16$ in Italian and $16$ in English), with the best-performing Italian prompt ultimately selected. Performance for VSV is measured with accuracy, i.e., the proportion of correctly selected true statements over the total statement pairs.
@@ 4.2 Task2: Open-ended VQA
OEVQA is a generative task, where models are tested on their ability to provide correct open-ended answers to a question related to video content. The model receives as input a prompt question and a video, and is tasked with generating a correct answer. The prompt used in the experiments was selected as the best-performing among $10$ tested variants ($5$ in Italian and $5$ in English). Generated responses are then evaluated according to two approaches: similarity-based and LLM as a judge.
Similarity-based metrics
compare a response against the pool of 8 reference answers available in MAIA. We used five token-level metrics: ROUGE Lin (2004), BLEU Papineni et al. (2002), BERT-Score Zhang et al. (2020), METEOR Lavie and Agarwal (2007) and CIDEr Vedantam et al. (2015).
LLM as a judge.
@@
Figure 3: Qwen2.5-VL’s fingerprint through MAIA’s reasoning categories: (a) illustrates model performance in NLU, Task 1, when TS-FS pairs are independent, while (b) reports performance on the same task with TS-FS pairs aggregated within their 8-items pool, penalizing models’ inconsistency; (c) visualizes the performance on NLG, Task 2.
@@
Black video InternVL2 0.18 0.07 0.40 0.43 0.08 0.06 0.73 0.03 0.03 0.03 0.02 0.20 0.05
Llava-Next-Video 0.00 0.00 0.00 0.01 0.00 0.00 0.00 0.01 0.01 0.00 0.00 0.01 0.01
Llava-oneVision 0.14 0.01 0.04 0.59 0.04 0.00 0.83 0.04 0.04 0.01 0.01 0.06 0.01
Qwen-2.5-VL 0.22 0.11 0.36 0.41 0.17 0.16 0.87 0.04 0.06 0.09 0.07 0.15 0.16
1-Frame InternVL2 0.28 0.32 0.41 0.32 0.11 0.33 0.47 0.17 0.33 0.15 0.39 0.11 0.21
Llava-Next-Video 0.08 0.12 0.23 0.01 0.03 0.21 0.01 0.06 0.11 0.03 0.14 0.01 0.03
Llava-oneVision 0.32 0.35 0.29 0.64 0.14 0.32 0.65 0.21 0.35 0.12 0.36 0.15 0.20
Qwen-2.5-VL 0.36 0.36 0.28 0.64 0.16 0.44 0.81 0.25 0.39 0.14 0.39 0.21 0.24
32-Frames InternVL2 0.31 0.41 0.35 0.38 0.12 0.42 0.39 0.28 0.38 0.18 0.43 0.11 0.30
Llava-Next-Video 0.03 0.04 0.04 0.01 0.03 0.09 0.00 0.01 0.06 0.01 0.04 0.01 0.01
Llava-oneVision 0.38 0.51 0.21 0.61 0.19 0.51 0.45 0.39 0.47 0.26 0.48 0.11 0.33
Qwen-2.5-VL 0.44 0.53 0.29 0.63 0.23 0.50 0.67 0.43 0.56 0.28 0.55 0.28 0.32
Table 1: VSV (Task $1$): aggregated accuracy of correct pools (8/8) across reasoning categories, penalizing models’ inconsistency.

### v3

@@ Abstract
We introduce MAIA (Multimodal AI Assessment), a native-Italian benchmark designed for fine-grained investigation of the reasoning abilities of visual language models on videos. MAIA differs from other available video benchmarks for its design, its reasoning categories, the metric it uses, and the language and culture of the videos. MAIA evaluates Vision Language Models (VLMs) on two aligned tasks: a visual statement verification task, and an open-ended visual question-answering task, both on the same set of video-related questions. It considers twelve reasoning categories that aim to disentangle language and vision relations by highlighting the role of the visual input. Thanks to its carefully taught design, it evaluates VLMs’ consistency and visually grounded natural language comprehension and generation simultaneously through an aggregated metric revealing low results that highlight models’ fragility. Last but not least, the video collection has been carefully selected to reflect the Italian culture, and the language data are produced by native-speakers. [^1] Data available at GitHub.
@@
Davide Testa1,2, Giovanni Bonetta2, Raffaella Bernardi3, Alessandro Bondielli4,5, Alessandro Lenci5, Alessio Miaschi6, Lucia Passaro4, Bernardo Magnini2 1Università di Roma La Sapienza, 2Fondazione Bruno Kessler (FBK), 3Free University of Bozen-Bolzano 4Dept. of Computer Science, University of Pisa, 5CoLing Lab, Dept. of Philology, Literature and Linguistics, University of Pisa 6Istituto di Linguistica Computazionale "A. Zampolli" (CNR-ILC), ItaliaNLP Lab, Pisa {dtesta, gbonetta, magnini}@fbk.eu, raffaella.bernardi@unibz.it {alessandro.bondielli, alessandro.lenci, lucia.passaro}@unipi.it, alessio.miaschi@ilc.cnr.it
@@ 1 Introduction
Figure 1: Structure of the MAIA benchmark. For each video there are two questions related to 12 reasoning categories. For each question there is a pool of 8 answers, each forming a question–answer pair associated with its own True–False statement pair. The example reported in the Figure is related to Out-of-scope reasoning category.
@@
To address these limitations, we present MAIA: a competence-oriented benchmark consisting of two paired tasks – multiple-choice Visual Statement Verification (VSV) and Open-Ended Visual Question Answering (OEVQA) – containing aligned datapoints. For instance, in the example in Figure 1, for the video showing a pizza cooking in a wood-fired oven, the VSV contains the True and False statements (TS and TF) A: There is no cake in the video and B: There is a cake in the video, and the OEVQA contains the (aligned) question What is the cake made of?. Such interleaved data let us evaluate models’ robustness: we evaluate a model positively, only if it performs correctly both on the visual statement verification task (NLU) – it selects There is no cake in the video – and on the Open-ended VQA (NLG) by generating something like There is a pizza, not a cake; while we evaluate it negatively, if it performs correctly only on one of the two tasks. Moreover, the VSV is organized in pools containing 8 TS-FS pairs that differ only on the surface, letting us evaluate the model’s consistency. We categorize questions based on the reasoning they elicit; for instance, the question in Figure 1 is Out-of-Scope. MAIA spans twelve reasoning categories, helping highlight the role of language and visual modalities across them. Finally, it implements an all-in-one evaluation philosophy that lets us evaluate both models’ robustness and consistency through its Aggregate Metric. Last but not least, MAIA is based on native Italian videos with language data obtained through human annotations and complemented by semi-automatic data augmentation. To the best of our knowledge, this is the first benchmark for the Italian language on videos.
@@ 2 Related Work
Various types of benchmarks have been proposed since the raise of VLMs. From the single task-oriented benchmarks (e.g., Antol et al. (2015); Das et al. (2017); Croce et al. (2021)), attention has now moved to task collections Xu et al. (2024); Lee et al. (2024b) in which models show impressive performance. As in the early phase Johnson et al. (2017); Shekhar et al. (2017); Suhr et al. (2017), such success is mitigated by the use of diagnostic benchmarks, such as Parcalabescu et al. (2022); Thrush et al. (2022); Chen et al. (2023); Bugliarello et al. (2023); Bianchi et al. (2024) and carefully curated benchmarks such as Xiao et al. (2024); Tong et al. (2024). The third type of benchmarks available focus on the VLMs competence in a holistic fashion, evaluating advanced perception and reasoning with domain-specific knowledge Yue et al. (2024b). A similar picture emerges for video-based VLMs. Here as well, early surveys call for careful evaluation (e.g., Zhong et al. (2022)), task-oriented benchmarks show impressive performance Grunde-McLaughlin et al. (2021); Yu et al. (2023), while fine-grained ones pinpoint important weaknesses Kesen et al. (2023), and competence-based analysis highlight there is significant room for improvement in multimodal video understanding (Patraucean et al., 2023). Finally, both Tong et al. (2024) for images and Kesen et al. (2023) for videos manage to highlight VLMs shortcomings by imposing a more stringent task-accuracy metric that account for model consistency across very similar data or correlated competencies. Thanks to the richness of MAIA data collection, we adopt such severe, and hence robust, evaluation code and propose a novel aggregate metric. Building on prior work, MAIA targets a low-resource language and the underexplored video domain. While some benchmarks Das et al. (2024); Zhang et al. (2023) include limited Italian multiple-choice tasks, none focus on high-level reasoning in Italian video contexts or analyze distributional biases in VLMs.
@@
Widely used benchmarks, such as AGQA Grunde-McLaughlin et al. (2021) and MVBench Li et al. (2024c) focus on explicit visual elements (e.g., entity, action, and the spatio-temporal reasoning involving them), instead MAIA’s categories focus on the interplay between language and vision, especially when this relation is implicit or must be inferred, a dimension largely neglected in prior Video QA benchmarks. Yue et al. (2024a) includes multiple-choice and open-ended data points from entirely independent data sets with different origins and content, and reports the average performance across distinct tasks. Instead, MAIA’s NLU and NLG data points are aligned, a unique feature of MAIA framework, as such it introduce an Aggregate Accuracy metric specifically designed to ensure that the performance of the model is evaluated consistently across multiple-choice and open-ended questions derived from the same underlying data. There are few other video-text benchmarks including both these formats (e.g., Zhou et al. (2025); Peng et al. (2024)), however, to the best of our knowledge, none of them attempt to define a unifying metric, as we do in MAIA.
@@ 3 The MAIA Benchmark
Figure 2: Overview of MAIA reasoning categories. For each of the 100 videos, it contains 2 questions for each of the 12 categories; for each question, it has 8 answers, and each of these answers has a corresponding TS-FS pair.
@@ 3.1 Dataset
We outline here the steps involved in creating the MAIA dataset while a comprehensive description of its construction, characteristics, as well as the validation and revision procedures can be found in Testa et al. (2025). Validation steps consists in a qualitative analysis and revision of the data, when necessary. [^2] Examples here are in English for readability.
@@
Questions and Answers Collection. The annotation process was carried out in two phases. In the first phase (question creation), $12$ qualified annotators wrote $2$ open-ended questions [^3] Yes/No and audio-based questions were prohibited. per video for each category, ensuring diversity in entities and events. [^4] Annotators were paid €100 for their work. A manual review verified adherence to guidelines and semantic categories. In the second phase (answer collection), we used Prolific [^5] https://www.prolific.com/ to solve the task, targeting Italian-native participants with specific cultural criteria (aged 25258080, born and raised in Italy, and Italian native speakers). Each annotator answered 12 out of 24 questions per video [^6] Annotators were paid £7 per hour., focusing on detailed, visually grounded responses. Each question was answered by eight annotators to guarantee both accuracy and variability within the pool.  [^7] This choice is supported by Mañas et al. (2024), who show that up to 8 demonstrations provide a good balance between diversity, accuracy, and efficiency when talking about in-context learning with LLMs for VQA evaluation. Two semi-automatic validation checks were applied to the collected answers: ($1$) semantic consistency with the corresponding question, and ($2$) contradiction tests across answers in the same pool. [^8] We found that 90.25% of the 8-answer pools exhibit full agreement, as they do not contain any contradictions. The remaining 9.75% was manually reviewed by an additional annotator to resolve inconsistencies. This shows agreement among annotators, that cannot be measured through standard annotator agreement metrics (e.g., Inter-Annotator Agreement) due to the open-ended nature of the task. After validation, the dataset consists of 2,4002{,}400 questions, each paired with a pool of $8$ high-quality answers, for a total of 19,20019{,}200 validated responses. Through a post-processing of the lexicon, we made sure that the final 8-answer pools are lexical diverse. [^9] The lexical overlap within the 8-answers pool is 21.9021.90%, similarly to the overlap between pairs randomly extracted from each pool (22.2122.21%). In addition, the average Type-Token Ratio (TTR) for content words within each pool is $0.55$.
@@
Statement Collection. As shown in Figure 2, TSs are descriptive declarative sentences that accurately align with the visual content of videos. TSs describe videos from different semantic perspectives, according to MAIA semantic categories. TSs were generated using GPT-4o (prompt in Figure 4A of the Appendix): for each question, it is given the 8 human generated answers and it is prompted to produce 8 TSs by combining the content of the question with the one of the corresponding answer. Again, post-processing techniques ensured high lexical variability within each pool reducing lexical overlap. FSs are incorrect descriptions automatically generated by prompting GPT-4o (Figure 4B) and created by minimally modifying elements of a TS related to a reasoning category while maintaining the original sentence structure, thus forming minimal pairs. FSs were validated through two semi-automatic checks (GPT-4o): (1) a structural verification to ensure that each FS was a minimal but incorrect variation consistent with its semantic category, and (2) an NLI-based contradiction test to confirm that each FS contradicted its corresponding TS. This process produced $19,200$ high-quality FSs aligned with their corresponding TSs.
@@ 3.2 Reasoning Categories
We report the reasoning categories in MAIA. [^10] More details in Appendix A.1. Figure 2 provides examples of aligned question, answer, TS, and FS.
@@ 4.1 Task1: Visual Statement Verification
VSV is a multiple-choice task where a model is presented with a true-false statement pair related to a MAIA question (see section 3.1) for a given video, and has to select the true option. The two statements are randomly assigned to two labels, A and B, and the model is asked to generate only the label. We chose the prompt through an extensive evaluation of $32$ variants ($16$ in Italian and $16$ in English), with the best-performing Italian prompt ultimately selected. Performance for VSV is measured with accuracy, i.e., the proportion of correctly selected true statements over the total statement pairs.
@@ 4.2 Task2: Open-ended VQA
OEVQA is a generative task, where models are tested on their ability to provide correct open-ended answers to a question related to video content. The model receives as input a prompt question and a video, and is tasked with generating a correct answer. The prompt used in the experiments was selected as the best-performing among $10$ tested variants ($5$ in Italian and $5$ in English). Generated responses are then evaluated according to the following approaches.
Similarity-based metrics.
It compares a response against the pool of 8 reference answers available in MAIA. We used five token-level metrics: ROUGE Lin (2004), BLEU Papineni et al. (2002), BERT-Score Zhang et al. (2020), METEOR Lavie and Agarwal (2007) and CIDEr Vedantam et al. (2015).
LLM-as-a-judge.
@@
Figure 3: Fingerprint of Qwen2.5-VL 72B through MAIA’s reasoning categories: (a) illustrates model performance in NLU, Task 1, when TS-FS pairs are independent, while (b) reports performance on the same task when the model correctly identify all TS-FS pairs within each 8-item pool, thereby penalizing inconsistency; (c) visualizes the performance on NLG, Task 2.
@@
Black video InternVL2 8B 0.18 0.07 0.40 0.43 0.08 0.06 0.73 0.03 0.03 0.03 0.02 0.20 0.05
InternVL3 78B 0.30 0.35 0.44 0.04 0.28 0.41 0.92 0.10 0.17 0.08 0.12 0.29 0.40
Llava-Next-Video 7B 0.00 0.00 0.00 0.01 0.00 0.00 0.00 0.01 0.01 0.00 0.00 0.01 0.01
Llava-oneVision 7B 0.14 0.01 0.04 0.59 0.04 0.00 0.83 0.04 0.04 0.01 0.01 0.06 0.01
Qwen-2.5-VL 7B 0.22 0.11 0.36 0.41 0.17 0.16 0.87 0.04 0.06 0.09 0.07 0.15 0.16
Qwen-2.5-VL 72B 0.22 0.17 0.31 0.72 0.17 0.02 0.92 0.03 0.08 0.01 0.03 0.11 0.08
1-Frame InternVL2 8B 0.28 0.32 0.41 0.32 0.11 0.33 0.47 0.17 0.33 0.15 0.39 0.11 0.21
InternVL3 78B 0.42 0.53 0.50 0.46 0.37 0.47 0.69 0.31 0.53 0.23 0.44 0.20 0.36
Llava-Next-Video 7B 0.08 0.12 0.23 0.01 0.03 0.21 0.01 0.06 0.11 0.03 0.14 0.01 0.03
Llava-oneVision 7B 0.32 0.35 0.29 0.64 0.14 0.32 0.65 0.21 0.35 0.12 0.36 0.15 0.20
Qwen-2.5-VL 7B 0.36 0.36 0.28 0.64 0.16 0.44 0.81 0.25 0.39 0.14 0.39 0.21 0.24
Qwen-2.5-VL 72B 0.40 0.40 0.32 0.72 0.22 0.34 0.81 0.34 0.49 0.21 0.47 0.15 0.25
32-Frames InternVL2 8B 0.31 0.41 0.35 0.38 0.12 0.42 0.39 0.28 0.38 0.18 0.43 0.11 0.30
InternVL3 78B 0.25 0.44 0.42 0.14 0.21 0.22 0.20 0.24 0.36 0.17 0.20 0.06 0.26
Llava-Next-Video 7B 0.03 0.04 0.04 0.01 0.03 0.09 0.00 0.01 0.06 0.01 0.04 0.01 0.01
Llava-oneVision 7B 0.38 0.51 0.21 0.61 0.19 0.51 0.45 0.39 0.47 0.26 0.48 0.11 0.33
Qwen-2.5-VL 7B 0.44 0.53 0.29 0.63 0.23 0.50 0.67 0.43 0.56 0.28 0.55 0.28 0.32
Qwen-2.5-VL 72B 0.54 0.67 0.41 0.80 0.39 0.59 0.75 0.56 0.65 0.39 0.65 0.13 0.46
Table 1: VSV (Task 11): accuracy of correct pools (8/8) across reasoning categories, penalizing models’ inconsistency.